OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

can someone please help me I am trying to finish my online course with an a and am stuck on this problem

OpenStudy (anonymous):

@Jim766 please help

OpenStudy (jim766):

\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}\] sorry I screwed the 1st one up..lol does this look right?

OpenStudy (anonymous):

I think so I tried to figure it out but I have learning problems lol

OpenStudy (jim766):

lol...let see if we can figure it out. we need to simplify each term if we can inside the radicals, I am looking for factors that are perfect squares, so when I square root them I get a number

OpenStudy (jim766):

\[3\sqrt{2x}+x \sqrt{4*2*x}-5\sqrt{9*2*x}\] the perfect squares are 4 and 9

OpenStudy (jim766):

\[3\sqrt{2x}+2*x \sqrt{2x}-5*3\sqrt{2x}\] when I take the sq root of the 4 and 9, I get this

OpenStudy (jim766):

\[3\sqrt{2x}+2x \sqrt{2x}-15\sqrt{2x}\] now that all the sq rts are the same I can combine like terms

OpenStudy (jim766):

\[3 -15+2x \sqrt{2x}\] \[-12 + 2x \sqrt{2x}\]

OpenStudy (jim766):

now that is the correct answer, but they got \[-10x \sqrt{2x}\] any idea what they did wrong?

OpenStudy (jim766):

they took -12 + 2x and got -10, you cant do that because they are not like terms, you cant combine them. that is why the correct answer leaves them separated. Does that help any?

OpenStudy (anonymous):

yes thank-you so like can you write it in the way of the question I cant get it right

OpenStudy (jim766):

3sqrt(2x) +xsqrt(8x) - 5sqrt(18x) 3sqrt(2x) +2xsqrt(2x) - 15sqrt(2x) (-12+2x)sqrt(2x) this is simplified

OpenStudy (anonymous):

thank-you soooooooo much!!!!!!